The number of days a school is in session for a year is 15 more than the number of days of vacation. How many days is the school in session out of 365 days in a year?
Answer by macston(5194)

.
V=number of vacation days; S=number of school days=V+15 days
V+S=365 days Substitute for S.
V+(V+15 days)=365 days Subtract 15 days from each side.
2V=350 days Divide each side by 2
V=175 days There are 175 vacation days per year.
S=V+15 days=175 days+15 days=190 days
ANSWER: The school is in session 190 days a year.